# Estimators for multivariate allometric regression model

### v2

@@ 5 Consistency under large $p$, large $n$ regimes
In this section, we consider “large $p$, large $n$” regimes under which $n$ and $p$ tend to positive infinity simultaneously. These regimes are frequently discussed in the literature. We suppose the following assumption.
@@
Under Assumption 2, $\beta>0$ implies that the largest eigenvalue is (asymptotically) spiked, while non-spiked eigenvalues imply $\beta\leq 0$. The cases of $\beta<0$ correspond to weak identifiability regimes. Moreover, $\delta<1$ implies $p/n\to\infty$.
@@
For $10<w<$ 1, it holds that
@@
$\leq\frac{8\{q(1-2w)+(n-1)w^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w)^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w\}(\lambda_{1}-\lambda_{2})+(1-w)\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(1)-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8\{n-1-q\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}}{(n-1-q)^{2}(\lambda_{1}-\lambda_{2})^{2}}=\frac{\Theta(p^{2+\delta})}{\Theta(p^{2\beta+2\delta})}=\Theta\left(p^{2-2\beta-\delta}\right),$
@@
$\leq\frac{8q\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left\{q(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right\}^{2}}$
$=\frac{\Theta(p^{2})+\Theta(p^{1+\delta})}{\left(\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\max\{2,1+\delta\}})}{\Theta(p^{\max\{2\beta,2\delta\}})}=\left\{\begin{array}[]{ll}\Theta\left(p^{1-\delta}\right)&(\delta\geq 1),\\ \Theta\left(p^{\min\{2-2\beta,2-2\delta\}}\right)&(\delta<1).\end{array}\right.$
@@
Since
$w^{*}$ =$=\frac{\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}}{2\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}(\lambda_{1}-\lambda_{2})q+2(\lambda_{1}+{\rm tr}(\bm{\Sigma}))(\lambda_{1}-\lambda_{2})\|\bm{X}\bm{\alpha}\|^{2}+\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}\|\bm{X}\bm{\alpha}\|^{2}}$
@@
we notice that, for $\delta\leq 1$, $w^{*}=\Theta(1)$ under $\beta\geq\delta$ and $w^{*}=\Theta(p^{\beta-\delta})$ under $\beta<\delta$. For $\delta>1$, $w^{*}=\Theta(p^{\beta-1})$. From $30\leq w^{*}<2/$ 3, it follows that $1-w^{*}=\Theta(1)$. We also note that
$\frac{8\{q(1-2w^{*})+(n-1)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}.$
@@
$\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=\frac{\Theta(p^{\max\{0,\delta+2\beta-2\}})\Theta(p^{2})+\Theta(p^{\delta+1})}{\left(\Theta(p^{\max\{0,\delta+\beta-1\}})\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\frac{\Theta(p^{\max\{2,\delta+1,\delta+2\beta\}})}{\left(\Theta(p^{\max\{\delta,\beta,\delta+2\beta-1\}})\right)^{2}}$
@@
$\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
@@
$\frac{8\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+16(1-w^{*})^{2}(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left(\{q(1-w^{*})+(n-1-q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right)^{2}}$
@@
We see that, as for sufficient conditions for the consistency, $\hat{\bm{\gamma}}_{1}(w^{*})$ has the good points both of $\hat{\bm{\gamma}}_{1}(w)$ with $10<w<$ 1 and $\hat{\bm{\gamma}}_{1}(0)$. This may be because the inequality $\|\cdot\|_{\mathrm{op}}\leq\|\cdot\|_{\mathrm{F}}$ is used. We may relax sufficient conditions in Theorem 5.1 by avoiding this inequality.
@@ 6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
$\hat{w}^{*}=\frac{q\{\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2}\}(\hat{\lambda}_{1}-\hat{\lambda}_{2})+2(\hat{\lambda}_{1}+{\rm tr}(\hat{\bm{\Sigma}}))(\hat{\lambda}_{1}-\hat{\lambda}_{2})\widehat{\|\bm{X}\bm{\alpha}\|^{2}}}{2q\{\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2}\}(\hat{\lambda}_{1}-\hat{\lambda}_{2})+2(\hat{\lambda}_{1}+{\rm tr}(\hat{\bm{\Sigma}}))(\hat{\lambda}_{1}-\hat{\lambda}_{2})\widehat{\|\bm{X}\bm{\alpha}\|^{2}}+\{\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2}\}\widehat{\|\bm{X}\bm{\alpha}\|^{2}}},$
@@
$\hat{\bm{\Sigma}}=\frac{\bm{S}_{E}}{n-1-q},$
$\widehat{{\rm tr}(\bm{\Sigma}^{2})}=\frac{1}{(n+1-q)(n-2-q)}\left\{{\rm tr}(\bm{S}_{E}^{2})-\frac{\left({\rm tr}(\bm{S}_{E})\right)^{2}}{n-1-q}\right\},$
$\widehat{\|\bm{X}\bm{\alpha}\|^{2}}={\rm tr}(\bm{S}_{R})-\frac{q}{n-1-q}{\rm tr}(\bm{S}_{E}),$
and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are the largest and second largest eigenvalues of $\hat{\bm{\Sigma}}$, respectively. It holds that $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ under some asymptotic regimes such as a traditional large sample regime where $n\to\infty$ with the assumption that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix while other parameters are fixed. Neither the consistency of $\hat{\bm{\gamma}}_{1}(w^{*})$ nor $\hat{w}^{*}$ necessarily means that a feasible estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is consistent up to sign, that is to say,
@@
Hence, let us discuss asymptotic regimes under which $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ implies (6.1).
@@
$\frac{\{q(1-2w^{*})+(n-1)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}\to 0$ ( 6.2)
@@
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}=O(1),$ ( 6.3)
if $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$, then (6.1) holds.
@@
Since $\bm{\gamma}_{1}$ is the first principal eigenvector of $\mathsf{E}[\bm{S}(w^{*})]$ and since
@@
$\sin\angle(\bm{\gamma}_{1},\hat{\bm{\gamma}}_{1}(\hat{w}^{*}))\leq\frac{2\|\bm{S}(\hat{w}^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{{\rm op}}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}\leq 2|\hat{w}^{*}-w^{*}|\frac{\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}+2\frac{\|\bm{S}(w^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{\rm op}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}.$
@@
because $\hat{w}^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}w^{*}$ is assumed. Since the Jensen inequality provides $\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}]\leq\sqrt{\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}^{2}]}$, we evaluate $\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|^{2}_{\rm op}]$. It follows from
$\|\bm{S}_{R}-\bm{S}_{E}\|_{\mathrm{op}}^{2}$ $\leq\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm F}^{2}={\rm tr}\Bigl{(}(\bm{S}_{R}-\bm{S}_{E})(\bm{S}_{R}-\bm{S}_{E})^{\top}\Bigr{)}$
@@
$\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\mathrm{op}}^{2}]\leq\mathsf{E}\left[{\rm tr}\left((\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])^{2}\right)\right]+\mathsf{E}\left[{\rm tr}\left((\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])^{2}\right)\right]+{\rm tr}\left((\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\right).$
@@
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-E[\bm{S}_{R}]\right)^{2}\right)\right]=q{\rm tr}(\bm{\Sigma}^{2})+q\left({\rm tr}(\bm{\Sigma})\right)^{2}+2\left(\lambda_{1}+{\rm tr}(\bm{\Sigma})\right)\|\bm{X}\bm{\alpha}\|^{2},$
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-E[\bm{S}_{E}]\right)^{2}\right)\right]=(n-1-q){\rm tr}(\bm{\Sigma}^{2})+(n-1-q)\left({\rm tr}(\bm{\Sigma})\right)^{2}.$
@@
${\rm tr}\left((\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\right)$ $={\rm tr}\left(\left\{-(n-1-2q)\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+\|\bm{X}\bm{\alpha}\|^{2}\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}\right\}^{2}\right)$
$=\left(n-1-2q\right)^{2}{\rm tr}(\bm{\Sigma}^{2})+\|\bm{X}\bm{\alpha}\|^{4}-2(n-1-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}.$
@@
$\leq\sqrt{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)({\rm tr}(\bm{\Sigma}))^{2}+\|\bm{X}\bm{\alpha}\|^{4}+2\{{\rm tr}(\bm{\Sigma})-(n-2-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}.$
Noting that $w^{*}<2/3$, $\lambda_{1}(w^{*})-\lambda_{2}(w^{*})=\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}$, and
$\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}+\|\bm{X}\bm{\alpha}\|^{4}+\{{\rm tr}(\bm{\Sigma})-(n-2-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
=$=\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}+\{{\rm tr}(\bm{\Sigma})-(n-2-2q)\lambda_{1}\}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}+O(1)$
=$=\frac{\{(n-1-2q)^{2}+n-1\}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma})\}^{2}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+(1-w^{*})\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}+o(1)+O(1),$
@@
The estimator $\hat{w}^{*}$ in Theorem 6.1 can be replaced with arbitrary consistent estimator of $w^{*}$.

### v4

@@ 5 Consistency under large $p$, large $n$ regimes
In this section, we consider “large $p$, large $n$” regimes under which $n$ and $p$ tend to positive infinity simultaneously. As introduced in Section 1, these regimes are frequently discussed in the literature. We suppose the following assumption.
@@
Under Assumption 2, $\beta>0$ implies that the largest eigenvalue is (asymptotically) spiked in the sense that it diverges to positive infinity, while non-spiked eigenvalues imply $\beta\leq 0$. The cases of $\beta<0$ correspond to weak identifiability regimes. Moreover, $\delta<1$ implies $p/n\to\infty$.
@@
In this proof, we use the notation (3.2) for simplicity and remark that $a=\Theta(p^{2})$, $b=\Theta(p)$, $c=\Theta(p^{\delta})$, $d=\Theta(p^{\beta})$. For $10<w<$ 1, it holds that
@@
$\leq\frac{8a\{q(1-2w)+(n-1)w^{2}\}+16bc(1-w)^{2}}{\left[d\{q+(n-1-2q)w\}+c(1-w)\right]^{2}}$
@@
$\mathsf{E}\left[\min_{\theta=-1,1}\|\theta\hat{\bm{\gamma}}_{1}(1)-\bm{\gamma}_{1}\|^{2}\right]\leq\frac{8a(n-1-q)}{d^{2}(n-1-q)^{2}}=\frac{\Theta(p^{2+\delta})}{\Theta(p^{2\beta+2\delta})}=\Theta\left(p^{2-2\beta-\delta}\right),$
@@
$\leq\frac{8aq+16bc}{(dq+c)^{2}}=\frac{\Theta(p^{2})+\Theta(p^{1+\delta})}{\left(\Theta(p^{\beta})+\Theta(p^{\delta})\right)^{2}}=\left\{\begin{array}[]{ll}\Theta\left(p^{1-\delta}\right)&(\delta\geq 1),\\ \Theta\left(p^{\min\{2-2\beta,2-2\delta\}}\right)&(\delta<1).\end{array}\right.$
@@
In this proof, we use the notation (3.2) for simplicity. Since
$w^{*}$ $=\frac{adq+2bcd}{2adq+2bcd+ac}$
@@
we notice that, for $\delta\leq 1$, $w^{*}=\Theta(1)$ under $\beta\geq\delta$ and $w^{*}=\Theta(p^{\beta-\delta})$ under $\beta<\delta$. For $\delta>1$, $w^{*}=\Theta(p^{\beta-1})$. From $30\leq w^{*}<2/$ 3, it follows that $1-w^{*}=\Theta(1)$. We also note that (3.5) is equal to
$\frac{8a\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}+16bc(1-w^{*})^{2}}{\left[d\{q(1-w^{*})+(n-1-q)w^{*}\}+c(1-w^{*})\right]^{2}}.$
@@
$\frac{8a\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}+16bc(1-w^{*})^{2}}{\left[d\{q(1-w^{*})+(n-1-q)w^{*}\}+c(1-w^{*})\right]^{2}}$
$=\frac{\Theta(p^{\max\{0,\delta+2\beta-2\}})\Theta(p^{2})+\Theta(p^{\delta+1})}{\left[\Theta(p^{\max\{0,\delta+\beta-1\}})\Theta(p^{\beta})+\Theta(p^{\delta})\right]^{2}}=\frac{\Theta(p^{\max\{2,\delta+1,\delta+2\beta\}})}{\left(\Theta(p^{\max\{\delta,\beta,\delta+2\beta-1\}})\right)^{2}}$
@@
$\frac{8a\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}+16bc(1-w^{*})^{2}}{\left[d\{q(1-w^{*})+(n-1-q)w^{*}\}+c(1-w^{*})\right]^{2}}$
@@
$\frac{8a\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}+16bc(1-w^{*})^{2}}{\left[d\{q(1-w^{*})+(n-1-q)w^{*}\}+c(1-w^{*})\right]^{2}}$
@@
We see that, as for sufficient conditions for the consistency, $\hat{\bm{\gamma}}_{1}(w^{*})$ has the good points both of $\hat{\bm{\gamma}}_{1}(w)$ with $10<w<$ 1 and $\hat{\bm{\gamma}}_{1}(0)$. We think that this may be because the inequality $\|\cdot\|_{\mathrm{op}}\leq\|\cdot\|_{\mathrm{F}}$ is used, so sufficient conditions in Theorem 5.1 may be relaxed by avoiding this inequality.
@@ 6 Feasible $\hat{\bm{\gamma}}_{1}(w^{*})$ estimator
$\hat{w}^{*}=\frac{\hat{a}\hat{d}q+2\hat{b}\hat{c}\hat{d}}{2\hat{a}\hat{d}q+2\hat{b}\hat{c}\hat{d}+\hat{a}\hat{c}}.$
@@
$\hat{a}=\widehat{{\rm tr}(\bm{\Sigma}^{2})}+({\rm tr}(\hat{\bm{\Sigma}}))^{2},\ \hat{b}=\hat{\lambda}_{1}+{\rm tr}(\hat{\bm{\Sigma}}),\ \hat{c}={\rm tr}(\bm{S}_{R})-q{\rm tr}(\hat{\bm{\Sigma}}),\ \hat{d}=\hat{\lambda}_{1}-\hat{\lambda}_{2},$
$\hat{\bm{\Sigma}}=\frac{\bm{S}_{E}}{n-1-q},\ \widehat{{\rm tr}(\bm{\Sigma}^{2})}=\frac{1}{(n+1-q)(n-2-q)}\left\{{\rm tr}(\bm{S}_{E}^{2})-\frac{\left({\rm tr}(\bm{S}_{E})\right)^{2}}{n-1-q}\right\},$
and $\hat{\lambda}_{1}$ and $\hat{\lambda}_{2}$ are the largest and second largest eigenvalues of $\hat{\bm{\Sigma}}$, respectively. When $\bm{\alpha}\neq\bm{0}_{q}$, noting that
$\mathsf{E}[{\rm tr}(\bm{S}_{R})]=q{\rm tr}(\bm{\Lambda})+\|\bm{X}\bm{\alpha}\|^{2},\quad\mathsf{V}[{\rm tr}(\bm{S}_{R})]=2q{\rm tr}(\bm{\Lambda}^{2})+4\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}$
and that $\bm{S}_{E}\sim\mathcal{W}_{p}(n-1-q,\bm{\Sigma})$, we can see that $\hat{w}^{*}-w^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ under some asymptotic regimes such as a traditional large sample regime where $n\to\infty$ with the assumption that $n^{-1}\bm{X}^{\prime}\bm{X}$ converges to a positive-definite matrix while other parameters are fixed. Neither the consistency of $\hat{\bm{\gamma}}_{1}(w^{*})$ nor $\hat{w}^{*}$ necessarily means that a feasible estimator $\hat{\bm{\gamma}}_{1}(\hat{w}^{*})$ is consistent up to sign, that is to say,
@@
Hence, let us discuss asymptotic regimes under which $\hat{w}^{*}-w^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ implies (6.1).
@@
$\frac{\{q(1-w^{*})^{2}+(n-1-q)(w^{*})^{2}\}\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}+(\lambda_{1}+{\rm tr}(\bm{\Sigma}))\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$\to 0$ ( 6.2)
@@
$\frac{(n-1-2q)^{2}{\rm tr}(\bm{\Sigma}^{2})+(n-1)\{{\rm tr}(\bm{\Sigma}^{2})+({\rm tr}(\bm{\Sigma}))^{2}\}-(n-2q)\lambda_{1}\|\bm{X}\bm{\alpha}\|^{2}}{\left[\{q+(n-1-2q)w^{*}\}(\lambda_{1}-\lambda_{2})+\|\bm{X}\bm{\alpha}\|^{2}\right]^{2}}$
$=O(1),$ ( 6.3)
if $\hat{w}^{*}-w^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$, then (6.1) holds.
@@
In this proof, we use the notation (3.2) for simplicity. Since $\bm{\gamma}_{1}$ is the first principal eigenvector of $\mathsf{E}[\bm{S}(w^{*})]$ and since
@@
$\sin\angle(\bm{\gamma}_{1},\hat{\bm{\gamma}}_{1}(\hat{w}^{*}))$ $\leq\frac{2\|\bm{S}(\hat{w}^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{{\rm op}}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}$
$\leq 2|\hat{w}^{*}-w^{*}|\frac{\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}+2\frac{\|\bm{S}(w^{*})-\mathsf{E}[\bm{S}(w^{*})]\|_{\rm op}}{\lambda_{1}(w^{*})-\lambda_{2}(w^{*})}.$
@@
because $\hat{w}^{*}-w^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$ is assumed. Since the Jensen inequality provides
$\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}]\leq\sqrt{\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm op}^{2}]},$
we evaluate $\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|^{2}_{\rm op}]$. It follows from
$\|\bm{S}_{R}-\bm{S}_{E}\|_{\mathrm{op}}^{2}$
$\leq\|\bm{S}_{R}-\bm{S}_{E}\|_{\rm F}^{2}={\rm tr}\Bigl{(}(\bm{S}_{R}-\bm{S}_{E})(\bm{S}_{R}-\bm{S}_{E})^{\top}\Bigr{)}$
@@
$\mathsf{E}[\|\bm{S}_{R}-\bm{S}_{E}\|_{\mathrm{op}}^{2}]$
$\leq\mathsf{E}\left[{\rm tr}\left((\bm{S}_{R}-\mathsf{E}[\bm{S}_{R}])^{2}\right)\right]+\mathsf{E}\left[{\rm tr}\left((\bm{S}_{E}-\mathsf{E}[\bm{S}_{E}])^{2}\right)\right]+{\rm tr}\left((\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\right).$
@@
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{R}-E[\bm{S}_{R}]\right)^{2}\right)\right]=aq+2bc$
and that
$\mathsf{E}\left[{\rm tr}\left(\left(\bm{S}_{E}-E[\bm{S}_{E}]\right)^{2}\right)\right]=a(n-1-q).$
@@
${\rm tr}\left((\mathsf{E}[\bm{S}_{R}-\bm{S}_{E}])^{2}\right)$ $={\rm tr}\left(\left\{-(n-1-2q)\bm{\Gamma}\bm{\Lambda}\bm{\Gamma}^{\prime}+c\bm{\gamma}_{1}\bm{\gamma}_{1}^{\prime}\right\}^{2}\right)$
$=\left(n-1-2q\right)^{2}{\rm tr}(\bm{\Sigma}^{2})+c^{2}-2c(n-1-2q)\lambda_{1}.$
@@
$\leq\sqrt{(n-1-2q)^{2}{\rm tr}(\bm{\Sigma}^{2})+a(n-1)+c^{2}+2c\{b-(n-1-2q)\lambda_{1}\}}.$
Noting that
$w^{*}<2/3,\ \lambda_{1}(w^{*})-\lambda_{2}(w^{*})=d\{q+(n-1-2q)w^{*}\}+c(1-w^{*}),$
and
$\frac{(n-1-2q)^{2}{\rm tr}(\bm{\Sigma}^{2})+a(n-1)+c^{2}+\{b-(n-1-2q)\lambda_{1}\}c}{\left[d\{q+(n-1-2q)w^{*}\}+c(1-w^{*})\right]^{2}}$
$=\frac{(n-1-2q)^{2}{\rm tr}(\bm{\Sigma}^{2})+a(n-1)+\{b-(n-1-2q)\lambda_{1}\}c}{\left[\{q+(n-1-2q)w^{*}\}d+c(1-w^{*})\right]^{2}}+O(1)$
$=\frac{(n-1-2q)^{2}{\rm tr}(\bm{\Sigma}^{2})+a(n-1)-c(n-2q)\lambda_{1}}{\left[\{q+(n-1-2q)w^{*}\}d+c(1-w^{*})\right]^{2}}+o(1)+O(1),$
@@
The estimator $\hat{w}^{*}$ in Theorem 6.1 can be replaced with arbitrary estimator $\hat{w}$ satisfying that $\hat{w}-w^{*}\stackrel{{\scriptstyle\mathrm{P}}}{{\to}}0$.